A1A is a proprietorship that has a calendar fiscal year. The proprietorship begins operations on April​ 1, 2020 and acquires a machine on December​ 1, 2020.

The machine has a cost of ​$22,000 and A1A incurs an additional ​$5,800 in expenses for installation. The machine is a Class 8 asset with a rate of 20​%.

What is the maximum CCA deduction A1A can take on this asset for the April 1 to December​ 31, 2020 fiscal​ year?

Choose the correct answer. ​(Round your answer to the nearest​ dollar.)

A. $4,189

B. $8,340

C. $6,284

D. $2,095
Answer #1

Calculation of CCA:

Total cost = Purchase Cost plus installation expenses = $(22000+5800) = $27800

Half year rule will be applied as this is first year of CCA deduction.

CCA = 27800* 1/2 * 20% * 275/365 = $ 2095

(Number of days are calculated from 1st April to 31st Dec, 2020 i.e. from the date of start of business operations = 275 days)

Correct answer is D. $ 2,095